Calculate the amount of heat released when one bottle (250g) of water is cooled from 30 degrees celcius to 25 degrees celciuis. The specific heat of water is (-4.186J/g degress celcius-).
1
Expert's answer
2020-04-19T15:30:03-0400

Solution:

To calculate the amount of heat entering or leaving a system, the equation Q = mcΔT is used.

m = mass (in grams)

c = specific heat capacity (J/g°C)

ΔT = change in temperature (°C).


ΔT = (Tfinal - Tinitial) = 25°C - 30°C = -5°C

Then,

Q = mcΔT = (250 g) * (4.186 J/g°C) * (-5°C) = -5232.5 J.

Q = -5232.5J.

The negative Joule value means heat is being lost in an exothermic reaction.


Answer: Q = -5232.5J.